how do I solve this question: (3.99-X)/3.5=1.35: how do I solve this question: (3.99-X)/3.5=1.35
Answer by sofiyacherni(96)

(3.99-X)/3.5=1.35/1
use the following formula if a/b=c/d then ad=bc
so (3.99-X)=3.5(1.35)
3.99-X=4.725
3.99-4.725=X
X=-0.735